# Interplay between local and global order controls bifurcations in nonreciprocal systems without a conservation law

### v1

@@ Abstract
We elucidate the generic bifurcation behavior of local and global order in short-range interacting nonreciprocal lattice systems without conservation laws. We show that a critical magnitude of nearest-neighbor correlations within the respective lattices controls the emergence of coherent oscillations of global order as a result of frustration. Local order is maintained during these oscillations, implying non-trivial spatio-temporal correlations. Ghost states emerge in the strong-interaction regime. The residence time in either of these states eventually diverges, giving rise to ordered non-equilibrium steady states via a saddle-node-infinite-period bifurcation. Our work provides a comprehensive microscopic understanding of collective phenomena in nonreciprocal matter without conservation laws, which may be relevant e.g. for understanding enzyme clustering.
The last decade saw a surge of interest in multicomponent systems with nonreciprocal interactions. Nonreciprocity is commonly found in active and non-equilibrium systems [1, 2, 3, 4, 5], such as non-equilibrium fluids [6], predator-prey [7, 8] and many-body lattice [9, 10, 11, 12, 13, 14] systems, spin glasses [15, 16, 17], active solids with odd elasticity [18], cell-cell interactions [19], optically trapped nanoparticles [20, 21], complex plasma dynamics [22, 23, 24], mixtures of passive and active particles [25] as well as chemically active interacting particles [26, 27, 28, 29], primitive metabolic cycles [30], quorum-sensing active particles [31], and non-Hermitian quantum systems [32, 33, 34]. The importance of nonreciprocal interactions is meanwhile also widely recognized in the social sciences [35, 36], neuroscience [37, 38], and robotics [39].
@@
An important aspect of nonreciprocal systems is the presence of conservation laws. In the case of two nonvariationally coupled fields one may have no [49, 9, 17, 50], one [51, 52, 53], or two [2, 25, 41, 42, 47, 44, 54] conservation laws. Nonreciprocal systems without conservation laws are particularly relevant in the realm of chemical reaction networks, where it was shown that kinetic asymmetries between enzymes may lead to nonreciprocal interactions [29].
There is growing interest in a bottom-up understanding of non-conserved nonreciprocal systems, so far limited to phenomenological and mean-field theory [49, 9, 17]. These works revealed intriguing phenomena, such as large-scale Hopf instabilities [49, 9, 17], saddle-node bifurcations [50, 49], and hidden collective oscillations [17]. However, a microscopic understanding of the bifurcation behavior remains elusive.
Here, we go beyond mean-field reasoning and develop a thermodynamically consistent description of nonreciprocal binary systems without conservation laws, starting from a pair of coupled Ising models. We incorporate nearest-neighbor correlations and consider both, global and local order. We explain why a critical magnitude of nearest-neighbor correlations controls the symmetry-breaking transition in the long-range order, in turn bounding the onset of coherent oscillations. We elucidate how the bifurcation behavior depends on inter-particle interactions, and highlight stark differences in the spatio-temporal dynamics of all-to-all (mean-field) versus short-range-interacting systems; the square and Bethe lattices display equivalent behavior that is strikingly different from the all-to-all lattice. Our work provides a deeper microscopic understanding of collective phenomena in nonreciprocal matter with more realistic short-ranged interactions.
@@
Figure 1: (a) Schematic of a pair of Ising lattices $a$ and $b$ with coordination number $z=4$ and cross-coupling $(K_{a},K_{b})$. For $K_{a}\neq K_{b}$ the cross-coupling is nonreciprocal. (b) Eigenvalues $\lambda_{\pm}$ of the linear stability matrix (see Eq. (8)) as a function of the steady state short-range order $q_{s}$ for $K_{a}=-K_{b}=0.1$. When $q_{s}\geq q_{\rm crit}\equiv 1/(z-1)$ (black vertical line) the real parts of $\lambda_{\pm}$ are non-negative, resulting in coherent oscillations. (c-f) Temporal evolution of the magnetization (c; see Eq. (2)), (d) short-range order (see Eqs. (3)-(4)), and (e,f) short-range correlations $\mathcal{C}^{\mu\nu}(t)\equiv q^{\mu\nu}(t)-m^{\mu}(t)m^{\nu}(t)$ for $K_{a}=-K_{b}=K=0.1$ and $J_{a}=J_{b}=0.4$. Black dashed lines in (c,d) are obtained with Monte-Carlo simulations on the Bethe lattice (see [56] for details). In all panels we consider $z=4$.
We consider single spin-flip Glauber dynamics [57]. Let $P(\boldsymbol{\sigma};t)$ be the probability at time $t$ to find the system in state $\boldsymbol{\sigma}=\{\sigma_{1}^{a},\sigma_{1}^{b},...,\sigma_{N}^{a},\sigma_{N}^{b}\}$ which is governed by the master equation $dP(\boldsymbol{\sigma};t)/dt=\sum_{\mu}\sum_{i}[w^{\mu}_{i}(-\sigma^{\mu}_{i})P(\boldsymbol{\sigma}^{\prime}_{\mu,i};t){-}w^{\mu}_{i}(\sigma^{\mu}_{i})P(\boldsymbol{\sigma};t)]$, where $\boldsymbol{\sigma}^{\prime}_{\mu,i}=\{\sigma_{1}^{a},\sigma_{1}^{b},...,-\sigma^{\mu}_{i},...,\sigma_{N}^{a},\sigma_{N}^{b}\}$ is a state which differs from state $\boldsymbol{\sigma}$ by one spin flip. The transition rates $w^{\mu}_{i}(\sigma^{\mu}_{i})$ to flip a spin $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$ are uniquely specified by limiting the interactions to nearest neighbors, imposing isotropy in position space, and requiring that for $K_{a}=K_{b}$ the transition rates obey detailed-balance. These physical restrictions then lead to the general result $w^{\mu}_{i}(\sigma^{\mu}_{i}){=}\left[1{-}\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\right]/2\tau$ [9, 49, 17], where $\Delta E^{\mu}_{i}=-2E^{\mu}_{i}$ is the change in energy on the $\mu=a\lor b$ lattice after spin conversion $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$, and $\tau\equiv 2N$ is an intrinsic timescale for spin flips to occur.
We are interested in the temporal dynamics of global and local order parameters averaged over all spins. The magnetization or long-range order [58] is given by
@@
where $\langle f(t)\rangle\equiv\sum_{\boldsymbol{\sigma}}P(\boldsymbol{\sigma};t)f(\boldsymbol{\sigma})$. Note that $m^{\mu}(t)\in[-1,1]$. The three short-range order parameters [58] are
@@
with correlations $\mathcal{C}^{\mu\nu}(t)\equiv q^{\mu\nu}(t)-m^{\mu}(t)m^{\nu}(t)$. We distinguish between the short-range order within and between the lattices. The alignment of spin pairs within lattice $\mu$ is quantified by $q^{\mu\mu}(t)$, and $q^{ab}(t)$ measures the alignment of equivalent spins on opposing lattices. The normalization in Eq. (3) follows from the fact that $zN/2$ is the number of nearest neighbor pairs in a periodic lattice with coordination number $z$, which renders $q^{\mu\mu}(t),q^{ab}(t)\in[-1,1]$.
@@
where $\sum_{l,n}\equiv\sum_{l=0}^{z}\sum_{n=0}^{1}$ is a sum over all possible nearest neighbor spin values on the same ($l=0,\ldots,z$) and opposing ($n=0,1$) lattice, $U^{\mu}_{l,n}\equiv[2l-z]J_{\mu}+[2n-1]K_{\mu}$, and $\mathcal{P}^{\mu}_{l,n}(t)\in[0,1]$ is the time-dependent probability to select an up or down spin which has $l$ neighboring up spins on the same lattice, and $n$ neighboring up spins on the opposing lattice. The probability is normalized as $\sum_{l,n}\mathcal{P}^{\mu}_{l,n}(t){=}1$. The equations (5) are not yet closed; evaluating $\mathcal{P}^{\mu}_{l,n}(t)$ for an arbitrary lattice is a daunting combinatorial task, since it depends on microscopic details and therefore on an infinite hierarchy of higher order parameters.
An accurate closed-form expression for $\mathcal{P}^{\mu}_{l,n}(t)$ can be obtained with the Bethe-Guggenheim (BG) approximation, where we assume perfect mixing of nearest neighbor spin pairs. Such an approximation is exact on loop-less lattices like the Bethe lattice [59], or large random graphs with fixed coordination number [60, 61]. As we show below, the spatio-temporal dynamics on these lattices agrees with the behavior on the square lattice. Note that a theory of the nonreciprocal Ising model on the fully-connected mean-field lattice has been constructed in [49], which we also report in Appendix A.
@@
where $\overline{\mathcal{P}}^{\pm}_{l}(q_{s})\equiv\mathcal{C}^{z}_{l}(1\mp q_{s})^{z-l}(1\pm q_{s})^{l}$ are the probabilities (6) evaluated at steady state values, and we introduced the critical short-range order $q_{\rm crit}\equiv 1/(z-1)$, which depends on the coordination number of the lattice. The solution of the linear stability equation can be written as $\mathbf{\delta m}(t)=\sum_{k=\pm}\mathcal{A}_{k}\mathrm{e}^{\lambda_{k}t/\tau}\boldsymbol{\nu}_{k}$, where $\mathcal{A}_{\pm}$ are set by the initial conditions, $\boldsymbol{\nu}_{\pm}$ are the eigenvectors of $\mathbf{M}(q_{s};J,K)$ and $\lambda_{\pm}(q_{s};J,K)$ the corresponding eigenvalues
@@
$\rm i$ being the imaginary unit. Since ${M_{ab}(q_{s};J,K\neq 0)\neq 0}$ (see proof in [56]), the eigenvalues are complex for $K\neq 0$ (see dashed lines in Fig. 1b), resulting in oscillatory perturbations. The Hopf bifurcation [62], also called type-${\rm II}_{\rm o}$ instability [63], occurs when complex conjugate eigenvalues transit the imaginary axis in the complex plane. According to Eq. (8) this occurs when $M_{aa}(q_{s};J,K)=0$ implying ${q_{s}(J,K)=q_{\rm crit}}$ as seen from Eq. (7). The Hopf bifurcation is thus set by the critical value $q_{\rm crit}$ for short-range order, and for $q_{s}>q_{\rm crit}$ we have ${\rm Re}(\lambda_{\pm})>0$ (solid line in Fig. 1b). In other words, when spins on the respective lattices are sufficiently aligned, a transition to an oscillatory state occurs (Fig. 1c,d). The critical value of the short-range order that determines the onset of coherent oscillations is our second main result that generalizes to other approximation schemes beyond the mean field approximation (see Appendix B).
@@
Figure 2: (a,b) Bifurcation diagram of the magnetization $(m^{a}_{s},m^{b}_{s})$ in the absence (a; $K=0$) and presence (b; $K=0.01$) of nonreciprocal coupling as a function of $J$. Inset of (b): Magnification of the bifurcation diagram around the Hopf bifurcation projected onto $m^{a}_{s}$. (c) Temporal evolution of the magnetization close to the degenerate saddle-node-infinite-period (SNIPER) bifurcation at $(J,K)=(0.35991,0.01)$. The magnetization oscillates coherently between $4$ ghost states ${\rm A}\rightarrow{\rm B}\rightarrow{\rm C}\rightarrow{\rm D}$ (for $K<0$ the direction is reversed), which eventually terminate in the $4$ respective stable branches at and beyond the SNIPER bifurcation. The life-time $\mathcal{T}$ in these ghost states diverges upon approaching the SNIPER bifurcation at $J_{\rm SNP}(K)$ from the left as $\lim_{J\uparrow J_{\rm SNP}(K)}\mathcal{T}\propto(J_{\rm SNP}(K)-J)^{-1/2}$. (d) Phase diagram of $m^{\mu}_{s}$: the gray region depicts the regime of coherent oscillations, with the black lines indicating isolines with fixed oscillation period, and the blue line indicating the Hopf bifurcations where $q_{s}=q_{\rm crit}\equiv 1/(z-1)$; in the light green region the magnetization is stationary and nonzero. In all panels we consider $z=4$.
Contrary to the idea that short-range order is slaved by long-range order [58], it rather seems that steady-state short-range order determines the bifurcation behavior of the long-range order. After sufficient short-range order is attained within the lattices, the frustration due to the nonreciprocal coupling gives rise to coherent oscillations: for $K>0$ a spin $\sigma^{a}_{i}$ wants to align with $\sigma^{b}_{i}$ that in turn tends to misalign with $\sigma^{a}_{i}$. This dynamical frustration results in an oscillatory motion of the order parameters [16]. Notably, for the one-dimensional lattice ($z=2$) we find $q_{\rm crit}=1$, which, in contrast to the mean-field prediction [49] (see also Appendix A), correctly implies the non-existence of a Hopf bifurcation.
@@
We start with the non-interacting case with $K=0$ (see Fig. 2a). For small values of $J$, there exists only one (trivial) stable steady state with $m^{\mu}_{s}=0$, as explained in the previous section. Increasing $J$ to $\ln{(z/(z-2))}/2$ we find a pitchfork bifurcation (red dot in Fig. 2a), which coincides with $q_{s}(J,0)=q_{\rm crit}$, and beyond which the trivial stable state becomes unstable. At the pitchfork bifurcation, $4$ stable branches (blue lines in Fig. 2a) and $4$ unstable branches (red dashed lines in Fig. 2a) emerge. The unstable branches are identified by a zero magnetization in either lattice, whereas the stable branches correspond to the spontaneously broken symmetry states with nonzero equilibrium magnetizations in both lattices.
Upon setting $K\neq 0$, the pitchfork bifurcation turns into a Hopf bifurcation (blue dot in Fig. 2b), whose $J$-value depends on $K$ through the relation $q_{s}(J,K)=q_{\rm crit}$ (blue line in Fig. 2d; see [56] for explicit results). Increasing $J$ beyond the Hopf bifurcation, there is a regime with coherent oscillations (gray area in Fig. 2b,d). In Fig. 2d we identify the isolines of fixed period of the oscillations (black lines). Upon further increasing $J$ at fixed $K$, we observe a non-linear transition from coherent oscillations to a non-zero stationary magnetization (light green region in Fig. 2d), which is set by a degenerate saddle-node-infinite-period (SNIPER) bifurcation (green dots in Fig. 2b). Approaching the SNIPER bifurcation from below, the magnetization in both lattices start to oscillate between four ghost states (see Fig. 2c), whereby the residence time within ghost states diverges as (see [56])
@@
where $J_{\rm SNP}(K)$ is the $J$-value of the SNIPER bifurcation at a given $K$ (green line in Fig. 2d). At the SNIPER bifurcation, one unstable and stable branch merge, resulting in $4$ unstable branches (red dashed lines in Fig. 2b) and $4$ stable branches (blue lines in Fig. 2b). Note that these stable states with non-zero stationary magnetization exist also in the square lattice system (see rigorous proof in [56]). This invalidates the hypothesis about the square lattice in Fig. 1d in [49].
The bifurcation diagram obtained with the mean-field approximation has similar qualitative features as in Fig. 2a,b (i.e, a Hopf and SNIPER bifurcation), however, the phase diagram displays a constant Hopf line at a fixed $J$, independent of $K$ (see [49] and Fig. A1). In Fig. 2d we see that the Hopf line with the BG approximation is $K$-dependent, closely resembling the empirical phase diagram on the qubic lattice (see [49]). This highlights the non-trivial role of short-range order that is not included in the mean-field approximation.
@@
Figure 3: (a-c) Spectral density $\langle|\hat{\Psi}^{a}_{k}(\omega)|\rangle$ of modes of coherent oscillations (Eq. (10)) on the (a) mean-field, (b) Bethe, and (c) square lattice. Note that the nonzero eigenvalues of the mean-field lattice are degenerate, which explains the white region in (a). The values for $J$ are chosen in the respective oscillatory regimes (see [56] for details), and $K=0.3$ for all three lattices. $\langle|\hat{\Psi}^{a}_{k}(\omega)|\rangle$ is averaged over $500$ independent trajectories, and $\omega_{0}={\rm argmax}\langle|\hat{\Psi}^{a}_{1}(\omega)|\rangle$ is the natural oscillation frequency. Resonances are visible at multiplies of $2\omega_{0}$, since we ignore the sign of projections. Insets: Temporal development of the projected microscopic states onto the eigenvectors of the Laplacian matrix of respective lattices. Colors denote three selected eigenvalues (and thus spatial scales) indicated by the dots in the main plot.
@@
Conclusion.—We explained the collective dynamics of the nonreciprocal Ising system on the level of both, local and global order. A critical threshold magnitude of nearest-neighbor correlations within the respective lattices was found to control the emergence of coherent oscillations of the global order parameter. Upon increasing interactions, ghost states emerge and the residence time in either of them eventually diverges, giving rise to a non-equilibrium steady state via a saddle-node-infinite-period bifurcation. Strikingly, during coherent oscillations of global order, a high degree of local order is preserved in nearest-neighbor-interacting systems. This implies non-trivial spatio-temporal correlations between spins, confirmed by a spectral-density maximum at large-scale low-frequency modes. In stark contrast, on the mean-field (all-to-all) lattice there is no distinction between different spatial modes, since each spin “feels” all spins, annihilating any notion of spatial structure. Thus, mean-field reasoning is insufficient; accounting for nearest-neighbor correlations is essential for a correct understanding of dynamics of nonreciprocal matter with short, or more generally finite, range of interactions.
Our work provides a comprehensive microscopic understanding of dynamic collective phenomena in nonreciprocal matter without conservation laws relevant, e.g. for enzyme clustering [68, 69]. What remains elusive are multiple ($>2$) coupled lattices, spatially heterogeneous/extended systems [70], as well as the thermodynamic cost of dynamical states and bifurcations [54, 71, 72, 73]. These will be addressed in future works.

### v2

@@ Abstract
We elucidate the generic bifurcation behavior of local and global order in the short-range interacting nonreciprocal Ising model evolving under Glauber dynamics. We show that a critical magnitude of nearest-neighbor correlations within the respective lattices controls the emergence of coherent oscillations of global order as a result of frustration. Local order is maintained during these oscillations, implying non-trivial spatio-temporal correlations. Long-lived states emerge in the strong-interaction regime. The residence time in either of these states eventually diverges, giving rise to ordered non-equilibrium steady states via a saddle-node-infinite-period bifurcation. Our work provides a comprehensive microscopic understanding of collective phenomena in nonreciprocal matter without conservation laws.
The last decade saw a surge of interest in multicomponent systems with nonreciprocal interactions. Nonreciprocity is commonly found in active and non-equilibrium systems [1, 2, 3, 4, 5], such as predator-prey [6, 7] and many-body lattice [8, 9, 10, 11, 12, 13] systems, spin glasses [14, 15, 16], active solids with odd elasticity [17], cell-cell interactions [18], optically trapped nanoparticles [19, 20], complex plasma dynamics [21, 22, 23], mixtures of passive and active particles [24] as well as chemically active interacting particles [25, 26, 27, 28, 29], primitive metabolic cycles [30], and non-Hermitian quantum systems [31, 32, 33]. The importance of nonreciprocal interactions is also widely recognized in the social sciences [34, 35], neuroscience [36, 37], and robotics [38]. However, the underlying mechanisms and implications of nonreciprocity can vary substantially among these disciplines.
@@
An important aspect of nonreciprocal systems is the presence of conservation laws. In the case of two nonvariationally coupled fields one may have no [48, 8, 16, 49], one [50, 51, 52], or two [2, 24, 40, 41, 46, 43, 53] conservation laws. Nonreciprocal systems without conservation laws are particularly relevant in the realm of chemical reaction networks, where it was shown that kinetic asymmetries between enzymes may lead to nonreciprocal interactions, resulting in spatio-temporal oscillations [28].
There is growing interest in a bottom-up understanding of non-conserved nonreciprocal systems, so far limited to phenomenological and mean-field theory [48, 8, 16]. These works revealed intriguing phenomena, such as large-scale Hopf instabilities [48, 8, 16], saddle-node bifurcations [49, 48], and hidden collective oscillations [16].
Here, we go beyond mean-field reasoning and develop a thermodynamically consistent description of nonreciprocal binary systems without conservation laws, starting from a pair of coupled Ising models, which serves as a minimal model of a many-body system with short-range interactions. We incorporate nearest-neighbor correlations and consider both, global and local order. We explain why a critical magnitude of nearest-neighbor correlations controls the symmetry-breaking transition in the global order, in turn bounding the onset of coherent oscillations. We elucidate how the bifurcation behavior depends on inter-particle interactions, and highlight stark differences in the spatio-temporal dynamics of all-to-all (mean-field) versus short-range-interacting systems; the square and Bethe lattices display equivalent behavior that is strikingly different from the all-to-all lattice.
@@
Figure 1: (a) Schematic of two Ising lattices $a$ and $b$ with coordination number $z{=}4$ and cross-coupling $(K_{a},K_{b})$. For $K_{a}\neq K_{b}$ the cross-coupling is nonreciprocal. (b) Eigenvalues $\lambda_{\pm}$ of the linear stability matrix (see Eq. (8)) as a function of the steady state local order $q_{s}$ for $K_{a}=-K_{b}=0.1$. When $q_{s}\geq q_{\rm crit}\equiv 1/(z-1)$ (black vertical line) the real parts of $\lambda_{\pm}$ are non-negative, resulting in coherent oscillations. (c-f) Temporal evolution of the magnetization (c; see Eq. (2)), (d) local order (see Eqs. (3)-(4)), and (e,f) local correlations $\mathcal{C}^{\mu\nu}(t)\equiv q^{\mu\nu}(t)-m^{\mu}(t)m^{\nu}(t)$ for $K_{a}=-K_{b}=K=0.1$ and $J_{a}=J_{b}=0.4$. Black dashed lines in (c,d) are obtained with Monte-Carlo simulations on the Bethe lattice (see [55] for details). In all panels we consider $z=4$.
We consider single spin-flip Glauber dynamics [56]. Let $P(\boldsymbol{\sigma};t)$ be the probability at time $t$ to find the system in state $\boldsymbol{\sigma}=\{\sigma_{1}^{a},\sigma_{1}^{b},...,\sigma_{N}^{a},\sigma_{N}^{b}\}$ which is governed by the master equation $dP(\boldsymbol{\sigma};t)/dt=\sum_{\mu}\sum_{i}[w^{\mu}_{i}(-\sigma^{\mu}_{i})P(\boldsymbol{\sigma}^{\prime}_{\mu,i};t){-}w^{\mu}_{i}(\sigma^{\mu}_{i})P(\boldsymbol{\sigma};t)]$, where $\boldsymbol{\sigma}^{\prime}_{\mu,i}=\{\sigma_{1}^{a},\sigma_{1}^{b},...,-\sigma^{\mu}_{i},...,\sigma_{N}^{a},\sigma_{N}^{b}\}$ is a state which differs from state $\boldsymbol{\sigma}$ by one spin flip. The transition rates $w^{\mu}_{i}(\sigma^{\mu}_{i})$ to flip a spin $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$ are uniquely specified by limiting the interactions to nearest neighbors, imposing isotropy in position space, and requiring that for $K_{a}=K_{b}$ the transition rates obey detailed-balance. These physical restrictions then lead to the general result $w^{\mu}_{i}(\sigma^{\mu}_{i}){=}\left[1{-}\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\right]/2\tau$ [8, 48, 16], where $\Delta E^{\mu}_{i}=-2E^{\mu}_{i}$ is the change in energy on the $\mu=a\lor b$ lattice after spin conversion $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$, and $\tau$ is an intrinsic timescale to attempt a single spin-flip.
We are interested in the temporal dynamics of global and local order parameters averaged over all spins. The magnetization or global order [57] is given by
@@
where $\langle f(t)\rangle\equiv\sum_{\boldsymbol{\sigma}}P(\boldsymbol{\sigma};t)f(\boldsymbol{\sigma})$. Note that $m^{\mu}(t)\in[-1,1]$. The three local order parameters [57] are
@@
with correlations $\mathcal{C}^{\mu\nu}(t)\equiv q^{\mu\nu}(t)-m^{\mu}(t)m^{\nu}(t)$. We distinguish between the local order within and between the lattices. The alignment of spin pairs within lattice $\mu$ is quantified by $q^{\mu\mu}(t)$, and $q^{ab}(t)$ measures the alignment of equivalent spins on opposing lattices. The normalization in Eq. (3) arises because $zN/2$ is the number of nearest neighbor pairs in a periodic lattice with coordination number $z$, which renders $q^{\mu\mu}(t),q^{ab}(t)\in[-1,1]$.
@@
where $\sum_{l,n}\equiv\sum_{l=0}^{z}\sum_{n=0}^{1}$ is a sum over all possible nearest neighbor spin values on the same ($l=0,\ldots,z$) and opposing ($n=0,1$) lattice, $U^{\mu}_{l,n}\equiv[2l-z]J_{\mu}+[2n-1]K_{\mu}$, and $\mathcal{P}^{\mu}_{l,n}(t)\in[0,1]$ is the time-dependent probability to select an up or down spin which has $l$ neighboring up spins on the same lattice, and $n$ neighboring up spins on the opposing lattice. The probability is normalized as $\sum_{l,n}\mathcal{P}^{\mu}_{l,n}(t){=}1$. Equations (5) are not yet closed; evaluating $\mathcal{P}^{\mu}_{l,n}(t)$ for an arbitrary lattice is a daunting combinatorial task, as it depends on microscopic details and therefore on an infinite hierarchy of order parameters.
An accurate closed-form expression for $\mathcal{P}^{\mu}_{l,n}(t)$ can be obtained with the Bethe-Guggenheim (BG) approximation, where we assume perfect mixing of nearest neighbor spin pairs. Such an approximation is exact on loop-less lattices like the Bethe lattice [58], or large random graphs with fixed coordination number [59, 60]. As we show below, the spatio-temporal dynamics on these lattices captures the behavior on the square lattice. Note that a theory of the nonreciprocal Ising model on the fully-connected mean-field lattice has been constructed in [48], which we also report in Appendix A.
@@
where $\overline{\mathcal{P}}^{\pm}_{l}(q_{s})\equiv\mathcal{C}^{z}_{l}(1\mp q_{s})^{z-l}(1\pm q_{s})^{l}$ are the probabilities (6) evaluated at steady state values, and we introduced the critical local order $q_{\rm crit}\equiv 1/(z-1)$, which depends on the coordination number of the lattice. The solution of the linear stability equation can be written as $\mathbf{\delta m}(t)=\sum_{k=\pm}\mathcal{A}_{k}\mathrm{e}^{\lambda_{k}t/\tau}\boldsymbol{\nu}_{k}$, where $\mathcal{A}_{\pm}$ are set by the initial conditions, $\boldsymbol{\nu}_{\pm}$ are the eigenvectors of $\mathbf{M}(q_{s};J,K)$ and $\lambda_{\pm}(q_{s};J,K)$ the corresponding eigenvalues
@@
$\rm i$ being the imaginary unit. Since ${M_{ab}(q_{s};J,K\neq 0)\neq 0}$ (see proof in [55]), the eigenvalues are complex for $K\neq 0$ (see dashed lines in Fig. 1b), resulting in oscillatory perturbations. The Hopf bifurcation [61], also called type-${\rm II}_{\rm o}$ instability [62], occurs when complex conjugate eigenvalues transit the imaginary axis in the complex plane. According to Eq. (8) this occurs when $M_{aa}(q_{s};J,K)=0$ implying ${q_{s}(J,K)=q_{\rm crit}}$ as seen from Eq. (7). The Hopf bifurcation is thus set by the critical value $q_{\rm crit}$ for local order, and for $q_{s}>q_{\rm crit}$ we have ${\rm Re}(\lambda_{\pm})>0$ (solid line in Fig. 1b). In other words, when spins on the respective lattices are sufficiently aligned, a transition to an oscillatory state occurs (Fig. 1c,d). The critical value of the local order that determines the onset of coherent oscillations is our second main result that generalizes to other approximation schemes beyond the mean field approximation (see Appendix B). After sufficient local order is attained within the lattices, the frustration due to the nonreciprocal coupling gives rise to coherent oscillations: for $K>0$ a spin $\sigma^{a}_{i}$ wants to align with $\sigma^{b}_{i}$ that in turn tends to misalign with $\sigma^{a}_{i}$. This dynamical frustration results in an oscillatory motion of the order parameters [15]. Notably, for the one-dimensional lattice ($z=2$) we find $q_{\rm crit}=1$, which, in contrast to the mean-field prediction [48] (see also Appendix A), correctly implies the non-existence of a Hopf bifurcation.
@@
Figure 2: (a,b) Bifurcation diagram of the magnetization $(m^{a}_{s},m^{b}_{s})$ in the absence (a; $K{=}0$) and presence (b; $K{=}0.01$) of nonreciprocal coupling as a function of $J$. Inset of (b): Magnification of the bifurcation diagram around the Hopf bifurcation projected onto $(m^{b}_{s},J)$-plane (see black box). (c) Temporal evolution of the magnetization close to the degenerate saddle-node-infinite-period (SNIPER) bifurcation at $(J,K){=}(0.35991,0.01)$. The magnetization oscillates coherently between $4$ ghost states ${\rm A}{\rightarrow}{\rm B}{\rightarrow}{\rm C}{\rightarrow}{\rm D}$ (for $K{<}0$ the direction is reversed), which eventually terminate in the $4$ respective stable branches at and beyond the SNIPER bifurcation. The life-time $\mathcal{T}$ in these ghost states diverges upon approaching the SNIPER bifurcation at $J_{\rm SNP}(K)$ from the left as $\lim_{J\uparrow J_{\rm SNP}(K)}\mathcal{T}\propto(J_{\rm SNP}(K)-J)^{-1/2}$. (d) Phase diagram of $m^{\mu}_{s}$: the gray region depicts the regime of coherent oscillations, with the black lines indicating isolines with fixed oscillation period, and the blue line indicating the Hopf bifurcations where $q_{s}{=}q_{\rm crit}{\equiv}1/(z{-}1)$; in the light green region the magnetization is stationary and nonzero. In all panels we consider $z=4$.
@@
We start with the non-interacting case with $K=0$ (see Fig. 2a). For small values of $J$, there exists only one (trivial) stable steady state with $m^{\mu}_{s}=0$, as explained in the previous section. Increasing $J$ to $\ln{(z/(z-2))}/2$ we find a pitchfork bifurcation (red dot in Fig. 2a), which coincides with $q_{s}(J,0)=q_{\rm crit}$, and beyond which the trivial stable state becomes unstable. At the pitchfork bifurcation, $4$ stable branches (blue lines in Fig. 2a) and $4$ unstable branches (red dashed lines in Fig. 2a) emerge. Unstable branches have zero magnetization in one of the lattices, while stable branches exhibit nonzero equilibrium magnetization due to spontaneously broken symmetry in both lattices.
Upon setting $K\neq 0$, the pitchfork bifurcation turns into a Hopf bifurcation (blue dot in Fig. 2b), whose $J$-value depends on $K$ through the relation $q_{s}(J,K)=q_{\rm crit}$ (blue line in Fig. 2d; see [55] for explicit results). Increasing $J$ beyond the Hopf bifurcation, there is a regime with coherent oscillations (gray area in Fig. 2b,d). In Fig. 2d we identify the isolines of fixed period of the oscillations (black lines). Upon further increasing $J$ at fixed $K$, we observe a non-linear transition from coherent oscillations to a non-zero stationary magnetization (light green region in Fig. 2d), which is set by a degenerate saddle-node-infinite-period (SNIPER) bifurcation (green dots in Fig. 2b). Approaching the SNIPER bifurcation from below, the magnetization in both lattices start to oscillate between four long-lived ghost states (see Fig. 2c). These $4$ long-lived states correspond to the virtual configuration, in which the respective lattices exert a “static” magnetic field on each other, and are characterized by a critical slowing down of the dynamics in the vicinity of the impending SNIPER bifurcation. The residence time within these ghost states diverges as (see [55])
@@
where $J_{\rm SNP}(K)$ is the $J$-value of the SNIPER bifurcation at a given $K$ (green line in Fig. 2d). At the SNIPER bifurcation, one unstable and stable branch emerge, resulting in $4$ unstable branches (red dashed lines in Fig. 2b) and $4$ stable branches (blue lines in Fig. 2b). Note that these stable states with non-zero stationary magnetization exist also in the finite square lattice system (see proof in [55]).
The bifurcation diagram obtained with the mean-field approximation has similar qualitative features as in Fig. 2a,b (i.e, a Hopf and SNIPER bifurcation), however, the phase diagram displays a constant Hopf line at a fixed $J$, independent of $K$ (see [48] and Fig. A1). In Fig. 2d we see that the Hopf line with the BG approximation is $K$-dependent, closely resembling the empirical phase diagram on the qubic lattice (see [48]).
@@
Figure 3: (a-c) Spectral density $\langle|\hat{\Psi}^{a}_{k}(\omega)|\rangle$ of modes of coherent oscillations (Eq. (10)) on the (a) mean-field, (b) Bethe, and (c) square lattice (see [55] for videos of the dynamics on the lattices). Note that the nonzero eigenvalues of the mean-field lattice are degenerate, which explains the white region in (a). The values for $J$ are chosen in the respective oscillatory regimes (see [55] for details), and $K=0.3$ for all three lattices. $\langle|\hat{\Psi}^{a}_{k}(\omega)|\rangle$ is averaged over $500$ independent trajectories, and $\omega_{0}={\rm argmax}\langle|\hat{\Psi}^{a}_{1}(\omega)|\rangle$ is the natural oscillation frequency. Resonances are visible at multiplies of $2\omega_{0}$, since we ignore the sign of projections. Insets: Temporal development of the projected microscopic states onto the eigenvectors of the Laplacian matrix of respective lattices. Colors denote three selected eigenvalues (and thus spatial scales) indicated by the dots in the main plot.
@@
Conclusion.—We explained the collective dynamics of the nonreciprocal Ising system on the level of both, local and global order. A critical threshold magnitude of local order within the respective lattices was found to control the emergence of coherent oscillations of the global order parameter. Upon increasing interactions, ghost states emerge and the residence time in either of them eventually diverges, giving rise to a non-equilibrium steady state via a saddle-node-infinite-period bifurcation. Strikingly, during coherent oscillations of global order, a high degree of local order is preserved in nearest-neighbor-interacting systems. This implies non-trivial spatio-temporal correlations between spins, confirmed by a spectral-density maximum at large-scale low-frequency modes. In stark contrast, on the mean-field (all-to-all) lattice there is no distinction between different spatial modes, annihilating any notion of spatial structure. Thus, accounting for nearest-neighbor correlations is essential for a correct understanding of dynamics of nonreciprocal matter with short, or more generally finite, range of interactions.
Our work provides a comprehensive microscopic understanding of dynamic collective phenomena in nonreciprocal matter without conservation laws relevant, e.g. for enzyme clustering [67, 68], where the nonreciprocal coupling relates to an attractive $K>0$ or repulsive $K<0$ interaction between two enzymes where the product of one is the substrate of the other enzyme. The magnetizations $m^{\mu}(t)$ in turn reflect time-dependent concentrations of product and substrates via $\phi^{\mu}(t)=(1+m^{\mu}(t))/2\in[0,1]$.
What remains elusive are multiple ($>2$) coupled lattices, spatially heterogeneous/extended systems [69], as well as the thermodynamic cost of dynamical states and bifurcations [53, 70, 71, 72]. Moreover, considering the relevance of conservation laws [73, 2, 41], it will be essential to develop a theoretical framework for the nonreciprocal Ising model with Kawasaki dynamics. These will be addressed in future works.